OpenStudy (anonymous):

combine like terms 7+4x-10x+15

OpenStudy (biokiiller396):

4x-10x= -6x 15+7=22 -6x+22

OpenStudy (kearsten):

the like terms are 4x-10x and 7+15

OpenStudy (anonymous):

sorry i forgot to put the answers